# Roadside Monocular 3D Detection Prompted by 2D Detection

### v3

@@ 3.1 The Proposed Framework: Pro3D
Table 1: Benchmarking results on the DAIR-V2X-I dataset. Following [56], we report AP metrics at IoU=0.5, 0.25, 0.25 in BEV on the three superclasses (Vehicle, Cyclist, and Pedestrian) respectively. We mark the modalties (M) used in each method: C for RGB camera, L for LiDAR. As our Pro3D can take in any 3D detectors, we build it with BEVDepth, BEVHeight, and BEVSpread, respectively. Results show that Pro3D significantly improves these methods, e.g., it helps BEVHeight get $>$10 AP improvements on Pedestrian and Cyclist!
@@
PointPillars [16] L 63.1 54.0 54.0 38.5 22.6 22.5 38.5 37.2 37.3
SECOND [54] L 71.5 54.0 54.0 54.7 31.1 31.2 55.2 52.5 52.5
MVXNet [42] L+C 71.0 53.7 53.8 54.1 30.8 31.1 55.8 54.5 54.4
FCOS3D [45] C 54.2 49.9 50.1 25.6 24.3 24.4 19.2 19.0 19.0
M3D-RPN [4] C 49.6 40.3 40.2 22.5 20.0 20.6 12.1 12.3 12.5
Kinematic3D [5] C 50.6 45.7 45.6 23.6 22.1 22.3 13.5 13.6 13.1
MonoDLE [28] C 52.3 48.2 48.1 24.1 23.5 23.6 14.6 14.4 14.5
MonoFlex [63] C 60.0 51.1 51.3 26.1 26.0 26.3 18.2 18.3 18.5
CaDDN [34] C 51.2 46.6 46.7 23.9 22.8 22.6 14.0 13.9 13.9
MonoNeRD [53] C 65.2 57.3 57.0 40.0 40.2 40.5 30.1 29.6 30.2
ImvoxelNet [37] C 44.8 37.6 37.6 21.1 13.6 13.2 6.8 6.7 6.7
BEVFormer [19] C 61.4 50.7 50.7 22.2 22.1 22.1 16.9 15.8 16.0
BEVFormer v2 [55] C 62.9 52.3 52.3 25.3 25.2 25.2 18.0 17.1 17.2
SparseBEV [21] C 65.6 55.1 52.6 30.0 30.3 30.0 18.4 17.9 17.5
MV2D [50] C 65.7 57.1 57.0 43.8 42.9 43.1 29.0 28.3 29.1
MonoGAE [57] C 84.6 75.9 74.2 44.0 47.6 46.8 25.7 24.3 24.4
BEVHeight++ [58] C 79.3 68.6 68.7 60.8 60.5 61.0 42.9 40.9 41.1
BEVDepth [18] C 75.5 63.6 63.7 55.7 55.5 55.3 35.0 33.4 33.3
@@
3.2 Discussions on 2D Detector Training
Our Pro3D exploits a 2D detector (Fig. 1a). We can train this 2D detector using roadside 3D detection datasets. As such datasets contain only 3D cuboid annotations, we derive 2D box annotations by using the camera extrinsics and intrinsics to project the 3D cuboids onto the 2D image plane. In principle, our method can use any 2D detectors, e.g., pretraining a powerful “base” 2D detector and finetuning on the target dataset. Below we make three remarks on this matter.
Techniques of 2D detection are well established. 2D detection is more sufficiently explored than monocular 3D detection, as the former has been a fundamental problem in computer vision whereas the latter is relatively new. There are many excellent 2D detectors, such as FasterRCNN [36], YOLO [3, 44], DETR [8] and DINO [62]. For benchmarking, we use the recent transformer-based detector DINO [62]. We also test other 2D detectors in experiments.
More datasets of 2D detection are publicly available. As 2D detection has been actively studied since the past decade, many large-scale datasets have been released to the public. In contrast, for the relatively new problem of monocular 3D detection, there are less datasets which are also small in scale compared with the 2D counterpart. For example, the 2D detection dataset COCO [20] (published in 2014) annotates 330K images with 2D boxes for 80 classes, whereas the 3D detection dataset nuScenes [7] (published in 2020) “only” annotates 144K images with 3D cuboids for 23 classes. These publicly available 2D datasets allows pretraining stronger 2D detectors, using which our Pro3D yields better 3D detection performance (Fig. 6).
Joint training vs. stage-wise training. For 3D detection, most methods jointly train the 3D detector and a 2D detector by incorporating 2D annotations in an auxiliary loss [18, 24, 30, 41, 46, 45, 63, 11, 55, 50, 27]. Our Pro3D is a stage-wise training method, which turns out to outperform joint training (Table 7). We conjecture this is due to two reasons. First, 2D detector training is better explored with mature techniques than 3D detector training, so stage-wise training makes use of such techniques of 2D detector training. Second, joint training must strike a balance between various loss functions such as 3D cuboid regression, rotation regression and so on; yet, optimizing over all losses inevitably sacrifices 2D detection which is actually much easier than 3D detection. It is worth noting that stage-wise training allows easy leverage of external 2D datasets for pretraining a stronger 2D detector, exploiting which helps our Pro3D achieve even better performance (Fig. 6).
Figure 5: Visual comparison between the state-of-the-art method BEVHeight [56] and our Pro3D. Results show that Pro3D can (1) make better orientation predictions than BEVHeight (column-1), and (2) detect objects (missed by BEVHeight) that are too small in size (column-2), heavily occluded (column-3), and in the far field (column-4).
Table 2: Benchmarking results on the Rope3D dataset. Following [59], we use metrics AP and Rope at IoU=0.5 and 0.7. Built with detectors BEVHeight and BEVSpread, Pro3D achieves significant improvements. Notably, it improves the original BEVHeight and BEVSpread by $>$26 AP on Big Vehicle! Conclusions in Table 1 hold on this dataset.
@@ 3.2 Discussions on 2D Detector Training
M3D-RPN [4] 54.2 62.7 33.1 44.9 16.8 32.9 6.9 24.2
Kinematic3D [5] 50.6 58.9 37.6 48.1 17.7 32.9 6.1 22.9
MonoDLE [28] 51.7 60.4 40.3 50.1 17.7 32.9 6.1 22.9
MonoFlex [63] 60.3 66.9 37.3 48.0 33.8 36.1 10.1 26.2
BEVFormer [19] 50.6 58.8 34.6 45.2 24.6 38.7 10.1 25.6
SparseBEV [21] 54.7 62.3 39.8 48.4 29.9 41.2 14.6 30.0
BEVDepth [18] 69.6 74.7 45.0 54.6 42.6 53.1 21.5 35.8
BEVHeight++ [58] 76.1 80.9 50.1 59.9 47.0 57.7 24.4 39.6
BEVHeight [56] 74.6 78.7 48.9 57.7 45.7 55.6 23.1 37.0
@@ 4 Experiments
We validate our Pro3D and design choices through extensive experiments. First, we present experiment settings including datasets, evaluation metrics and implementations details. Then, we compare our Pro3D against prior works. Lastly, we conduct comprehensive ablation studies. We refer the reader to the supplement for more analyses.
@@ 4.1 Settings
Datasets. We use two publicly available datasets in experiments: DAIR-V2X-I [61], and Rope3D [59].
  • DAIR-V2X-I is the first large-scale dataset supporting the research of roadside monocular 3D detection. It provides 5K and 2K images for training and validation, respectively. Following [56], we report results on the validation set, with breakdown analysis w.r.t three superclasses, “vehicle”, “cyclist”, and “pedestrian”, at three difficulty levels (Easy, Medium, Hard).
  • Rope3D is another large-scale dataset that contains 50k images and annotations of $>$1.5M objects in diverse scenes. Following [56], we use its train-set (70%) to train models and the validation set (30%) for benchmarking.
@@
  • Mean AP (mAP) [20] is used to report 2D detection performance on the image plane (Table 5 and Fig. 6). It first computes per-class average precision (AP) over IoU threshods from 0.5 to 0.95 with a step-size 0.05, and then returns the mean of per-class APs.
@@
Compared methods. Our framework Pro3D can take in any 2D detectors and 3D detectors. To show the benefits of Pro3D, we build it with representative roadside 3D detectors, including BEVDepth [18], BEVHeight [56], and BEVSpread [48]. For benchmarking, we compare against more methods which are either published recently or representative in the literature 3D detection. We refer the reader to Table 1 and 2 for the compared methods.
@@
Implementations. 3D detectors studied in our work mostly follow BEVHeight [56] and BEVSpread [48], which contain an image encoder with ResNet101 architecture [13], a BEV encoder with ResNet18 architecture, an FPN for proposal detection [54], and a detector head for the final output of 3D detections. It is worth noting that our Pro3D replaces the ResNet18-based BEV encoder with a single convolutional layer, removing the FPN as our Pro3D uses 2D detection already. This helps our Pro3D achieve faster inference than these previous mehtods (Table 9). Within our Pro3D framework, we report performance by using BEVDepth, BEVHeight and BEVSpread as the 3D detection module, respectively. As BEVHeight has the fastest inference speed (detailed in the supplement), we use it in our Pro3D for ablation study. For 2D detector, we train DINO [62] with the backbone ViT-S/16 as the 2D detector using the default hyperparameter setting. For all methods, the input image resolution is 864$\times$1536 and the initial voxel resolution is 1024$\times$1024$\times$1. We use random scaling and rotation augmentation during training, and train all models for 50 epochs (long enough for convergence) using AdamW optimizer [26] with learning rate 8e-4. We select the best checkpoint via validation. As a post-processing step, we remove detections that have confidence scores lower than 0.3. The supplementary material contains more detailed implementations, video demo, and code.
@@ 4.2 Benchmarking Results
We study how much improvements our Pro3D brings to existing roadside 3D detectors on DAIR-V2X-I (Table 1) and Rope3D (Table 2). Fig. 5 visualizes results by Pro3D and BEVHeight (refer to the supplement for more visual comparisons). The results demonstrate that our Pro3D helps previous methods BEVDepth [18], BEVHeight [56], and BEVSpread [48] achieve significant performance gains. Moreover, Pro3D boosts 3D detection performance particularly on small objects of “cyclist” and “pedestrian”.
@@ 4.3 Ablation Study and Analysis
Incorporating scene prior improves 3D detection. We concatenate a testing image and the scene prior (a “blank” RGB background image as shonw in Fig. 4) as a 6-channel input to the 3D detector. Table 3 shows that adding the scene prior remarkably improves roadside 3D detection. To study how different scene prior affects the final performance, we generate scene priors over training frames captured at night. Table 4 shows that our method is quite robust to scene priors that have different illuminations.
Figure 6: We train different 2D detectors (YOLOV7 [44] and DINO [62]) on different amounts of 2D annotations (provided by nuImages [7]) for Pro3D. We measure the 3D detection by averaging AP over the three classes vehicle, cyclist and pedestrian. Training on more 2D annotations leads to better 2D detectors and importantly, increases 3D detection performance of Pro3D.
Exploiting a stronger 2D detection yields better 3D detection performance. We compare methods’ performance w.r.t 2D detection metrics in Table 5. This helps us understand that (1) 2D detection performance greatly reflects the 3D detection performance; (2) exploiting a stronger 2D detector helps our Pro3D achieve better 3D detection performance. We train stronger 2D detectors (e.g., YOLOV7 [44] and DINO [62]) on more 2D annotations from nuImages [7]. Fig. 6 shows 3D detection performance by our Pro3D versus 2D detection performance (by training different 2D detectors). Results show that exploiting stronger 2D detectors greatly improves 3D detection performance with our Pro3D.
Table 5: Comparison w.r.t 2D detection metric (mAP) on DAIR-V2X-I. For both 2D detector (DINO) and 3D detectors BEVDepth and BEVHeight, we train them on the training set of DARI-V2X-I. To evaluate 3D detectors w.r.t 2D metrics, we project their 3D detections onto the 2D image plane to derive their predicted 2D boxes. Compared to BEVDepth and BEVHeight, DINO achieves $>$10 mAP higher on all classes and $\sim$20 mAP higher on Pedestrian! Owing to the effective leverage of the 2D detector DINO, our Pro3D significantly boosts performance of the compared methods.
@@
2D detector (DINO) [62] 74.3 75.6 75.6 57.3 54.7 54.9 50.6 51.6 51.6
BEVDepth [18] 62.9 64.8 64.2 41.5 44.8 44.9 32.9 33.1 33.0
Pro3D w/ BEVDepth 68.0 68.3 68.9 45.9 47.7 47.6 42.1 42.6 42.6
BEVHeight [56] 63.5 64.7 64.5 41.4 44.6 44.7 33.6 33.7 33.9
Pro3D w/ BEVHeight 68.1 68.8 68.8 47.9 49.5 49.4 42.8 43.5 43.5
BEVSpread [48] 64.1 65.2 64.6 41.7 45.1 44.9 33.9 34.1 34.3
Pro3D w/ BEVSpread 68.7 69.4 69.2 48.7 50.3 50.3 44.1 44.9 44.9
Table 6: Comparison of different prompting methods. As a baseline, we train BEVHeight with scene prior (Table 3). First, all prompting methods achieve remarkable improvements over BEVHeight. Moreover, using the center of predicted 2D box coordinates (“fuse 2D centers”) as prompts outperforms feature prompts. Lastly, additionally incorporating predicted labels brings further improvements.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
BEVHeight [56] 79.2 67.7 67.8 62.5 62.3 62.7 44.3 42.5 42.7
@@
What to use as prompts. To exploit the 2D detector to help train 3D detectors, we study what information from it can be used as prompts, including its intermediate feature map, feature vector located at a 2D detection, and even the simplistic 2D detection output (i.e., predicted box coordinates and class label). Note that they require different fusion methods to accommodate their different dimensions. For example, when using feature maps, one can concatenate them with feature of the encoder in the 3D detector, while using feature vectors or detections output require fusing vectors with the feature maps. Therefore, we use concatenation to fuse feature maps, and an attention module to fuse the feature vectors. Table 6 lists the results. Surprisingly, using the simple detections (2D boxes and class labels) as prompts performs better than feature maps and feature vectors located at 2D detections, where 2D boxes as prompts achieves [68.5, 66.2, 47.2] AP for Hard examples of the three classes [vehicle, cyclist, pedestrian]. Furthermore, additionally incorporating the predicted class labels in prompts performs the best, boosting to [68.9, 68.2, 50.1] AP on Hard examples of the three classes! In hindsight, we see that fusing box coordinates and labels provides both semantic and spatial information for the 3D detector to leverage, facilitating 3D detection. In contrast, fusing feature maps does not pinpoint the location of the objects and hence does not simplify 3D detector training.
Table 7: Comparison of stage-wise training (SW) vs. joint training (JT) for our Pro3D. We report AP metrics at IoU=0.5, 0.25, 0.25 in BEV. As a baseline, we train Pro3D with BEVHeight and scene prior. It is worth noting that stage-wise training allows us to train separate 2D and 3D detector backbones, i.e., their backbones are not shared (NS). Due to this, we can exploit more 2D annotated data (MD) to train better 2D detectors, e.g., using nuImages [7] to pretrain a 2D detector. Results convincingly show that stage-wise training outperforms joint training for roadside monocular 3D detection. With NS and MD, our Pro3D yields more notable improvements.
@@
Stage-wise training vs. joint training. We test different training paradigms in our Pro3D, e.g., (1) joint end-to-end training the 3D detector using both 2D and 3D losses, and (2) stage-wise training the 2D detector first and then the 3D detector. Note that stage-wise training allows us to train separate 2D and 3D detector backbones, i.e., their backbones are not shared (NS). This further makes it convenient to exploit more 2D annotated data (MD) to train better 2D detectors, e.g., using nuImages [7] to pretrain a 2D detector. Results in Table 7 show that stage-wise training outperforms joint training for roadside monocular 3D detection.
How to train 3D detectors with 2D detection prompts. The Pro3D uses 2D detection to prompt the 3D detector to make 3D predictions. To train the 3D detector, we can either use ground-truth 2D boxes or predicted ones as prompts. We compare their performance w.r.t AP in BEV on the DAIR-V2X-I benchmark. Results in Table 8 show that directly using the predicted 2D boxes (with predicted class labels) as prompts performs better.
@@
Pro3D w/ grnd 82.7 70.1 70.3 70.5 70.4 70.6 52.8 51.7 51.3
Pro3D w/ pred 82.9 70.7 70.8 71.3 71.2 71.4 53.9 52.9 52.6
Latency and parameters. Recall that our Pro3D achieves the best performance by using separate 2D and 3D detectors without sharing the visual encoder (Table 7). We study this model in Table 9 w.r.t model parameters, the wall-clock time, and FPS in processing a single RGB frame. Here, we run the 2D detector (either YOLOV7 or DINO) and the 3D detector encoder in parallel. Note that our Pro3D does not require an FPN for proposal detection because the 2D detector provides pinpointed detections already. Our Pro3D also replaces ResNet18 encoder as adopted in prior methods [56] with a single convolutional layer (Sec. 4.1). These changes make our Pro3D faster than the original BEVHeight in inference, regardless exploiting either YOLOV7 [44] or DINO [56] as the 2D detector. In summary, our Pro3D model “w/ YOLOV7” boosts 3D detection performance by 7.9 AP over the prior art BEVHeight, with faster inference speed (15.4 vs. 13.0 FPS) and fewer parameters (86.6M vs. 94.7M)!
Table 9: Comparison of inference time and parameters between our Pro3D and BEVHeight. AP means the average AP in BEV of all the three classes in DAIR-V2X-I; “#param” counts the parameters (in million); time is the wall-clock in ms, FPS is frames per second. We use two different 2D detectors in our Pro3D, namely YOLOV7 and DINO. We list the number of parameters of 3D detector plus 2D detector. Our model Pro3D “w/ YOLOV7” boosts 3D detection performance by 7.9 AP over the prior art BEVHeight, with faster inference speed (15.4 vs. 13.0 FPS) and fewer parameters (86.6M vs. 94.7M)!

### v4

@@ 3.2 Discussions on 2D Detector Training
Table 1: Benchmarking on DAIR-V2X-I. Following [yang2023BEVHeight], we report AP metrics at IoU=0.5, 0.25, 0.25 in BEV on Vehicle, Cyclist, and Pedestrian, respectively. We mark the modalties (M) used in each method: C for RGB camera, L for LiDAR. As our Pro3D can take in any 3D detectors, we build it with BEVDepth, BEVHeight, and BEVSpread, respectively. Results show that Pro3D significantly improves these methods, e.g., it helps BEVHeight achieve >>10 AP improvements on Pedestrian and Cyclist!
@@
PointPillars [lang2019pointpillars] L 63.1 54.0 54.0 38.5 22.6 22.5 38.5 37.2 37.3
SECOND [yan2018second] L 71.5 54.0 54.0 54.7 31.1 31.2 55.2 52.5 52.5
MVXNet [sindagi2019mvxnet] L+C 71.0 53.7 53.8 54.1 30.8 31.1 55.8 54.5 54.4
proj. w/ class prior C 33.2 24.6 24.5 15.2 14.9 14.8 13.5 13.3 13.4
proj. w/ learning C 39.2 29.9 29.7 18.6 18.3 18.3 13.9 13.8 13.9
FCOS3D [wang2021fcos3d] C 54.2 49.9 50.1 25.6 24.3 24.4 19.2 19.0 19.0
FCOS3D+ C 59.3 55.2 55.7 35.2 34.8 34.9 27.9 27.5 27.6
M3D-RPN [brazil2019m3d] C 49.6 40.3 40.2 22.5 20.0 20.6 12.1 12.3 12.5
Kinematic3D [brazil2020kinematic] C 50.6 45.7 45.6 23.6 22.1 22.3 13.5 13.6 13.1
MonoDLE [ma2021delving] C 52.3 48.2 48.1 24.1 23.5 23.6 14.6 14.4 14.5
MonoFlex [zhang2021objects] C 60.0 51.1 51.3 26.1 26.0 26.3 18.2 18.3 18.5
CaDDN [reading2021categorical] C 51.2 46.6 46.7 23.9 22.8 22.6 14.0 13.9 13.9
MonoNeRD [xu2023mononerd] C 65.2 57.3 57.0 40.0 40.2 40.5 30.1 29.6 30.2
ImvoxelNet [rukhovich2022imvoxelnet] C 44.8 37.6 37.6 21.1 13.6 13.2 6.8 6.7 6.7
BEVFormer [li2022bevformer] C 61.4 50.7 50.7 22.2 22.1 22.1 16.9 15.8 16.0
BEVFormer v2 [yang2023bevformer] C 62.9 52.3 52.3 25.3 25.2 25.2 18.0 17.1 17.2
SparseBEV [liu2023sparsebev] C 65.6 55.1 52.6 30.0 30.3 30.0 18.4 17.9 17.5
MV2D [wang2023object] C 65.7 57.1 57.0 43.8 42.9 43.1 29.0 28.3 29.1
MonoGAE [yang2024monogae] C 84.6 75.9 74.2 44.0 47.6 46.8 25.7 24.3 24.4
BEVHeight++ [yang2023bevheight++] C 79.3 68.6 68.7 60.8 60.5 61.0 42.9 40.9 41.1
BEVDepth [li2023BEVDepth] C 75.5 63.6 63.7 55.7 55.5 55.3 35.0 33.4 33.3
@@
Table 2: Benchmarking results on the Rope3D dataset. Following [ye2022rope3d], we use metrics AP and Rope at IoU=0.5 and 0.7. Built with detectors BEVHeight and BEVSpread, Pro3D achieves significant improvements. Notably, it improves the original BEVHeight and BEVSpread by >>26 AP on Big Vehicle! Conclusions in Table 1 hold on this dataset.
@@
M3D-RPN [brazil2019m3d] 54.2 62.7 33.1 44.9 16.8 32.9 6.9 24.2
Kinematic3D [brazil2020kinematic] 50.6 58.9 37.6 48.1 17.7 32.9 6.1 22.9
MonoDLE [ma2021delving] 51.7 60.4 40.3 50.1 17.7 32.9 6.1 22.9
MonoFlex [zhang2021objects] 60.3 66.9 37.3 48.0 33.8 36.1 10.1 26.2
BEVFormer [li2022bevformer] 50.6 58.8 34.6 45.2 24.6 38.7 10.1 25.6
SparseBEV [liu2023sparsebev] 54.7 62.3 39.8 48.4 29.9 41.2 14.6 30.0
BEVDepth [li2023BEVDepth] 69.6 74.7 45.0 54.6 42.6 53.1 21.5 35.8
BEVHeight++ [yang2023bevheight++] 76.1 80.9 50.1 59.9 47.0 57.7 24.4 39.6
BEVHeight [yang2023BEVHeight] 74.6 78.7 48.9 57.7 45.7 55.6 23.1 37.0
@@ 4 Experiments
We validate Pro3D through extensive experiments. We first describe experiment settings including datasets, metrics and implementation details. We then compare Pro3D against existing works. We lastly conduct comprehensive ablation study. The supplement provide more analyses.
@@ 4.1 Settings
Datasets. We use two established datasets:
  • DAIR-V2X-I [yu2022dair] is the first dataset of roadside monocular 3D detection. It provides 5K and 2K images for training and validation, respectively. Following [yang2023BEVHeight], we report results on the validation set, with breakdown analysis w.r.t three superclasses (vehicle, cyclist, and pedestrian), at three difficulty levels (Easy, Medium, Hard).
  • Rope3D [ye2022rope3d] is another large-scale dataset that contains 50K images and annotations of >>1.5M objects in diverse scenes. Following [yang2023BEVHeight], we use its train-set (70%) to train models and the validation set (30%) for benchmarking.
@@
  • Mean AP (mAP) [lin2014coco] measures 2D detection performance on the image plane (Fig. 7). It first computes per-class average precision (AP) over IoU thresholds from 0.5 to 0.95 with stepsize 0.05, and then averages per-class APs.
@@
Compared methods. Pro3D can take in any 2D detectors and 3D detectors. To show the benefits of Pro3D, we build it with representative roadside 3D detectors, including BEVDepth [li2023BEVDepth], BEVHeight [yang2023BEVHeight], and BEVSpread [wang2024bevspread]. For benchmarking, we compare against more methods which are either published recently or representative in the literature 3D detection. We refer the reader to Table 1 and 2 for the compared methods.
@@
[…]
@@
Figure 6: Comparison of 3D detections by Pro3D on a test image of the DAIR-V2X-I dataset. Ground-truth and predictions are green and red, respectively. The area in red boxes reveals a clear improvement in pedestrian (i.e., small objects) detection when the scene prior is used. Zoom in to see better.
@@ 4.2 Benchmarking Results
We study how much improvements our Pro3D brings to existing roadside 3D detectors on DAIR-V2X-I (Table 1) and Rope3D (Table 2). Fig. 5 visualizes results by Pro3D and BEVHeight (refer to the supplement for more visual comparisons). The results demonstrate that our Pro3D helps previous methods BEVDepth [li2023BEVDepth], BEVHeight [yang2023BEVHeight], and BEVSpread [wang2024bevspread] achieve significant performance gains. Importantly, Pro3D boosts 3D detection performance particularly on small objects of “cyclist” and “pedestrian”.
A “traditional” baseline is to use camera intrinsics and extrinsics to project 2D detections to 3D. It can roughly localize a detected 2D box in 3D but cannot determine the object’s pose and shape. We address this with two methods: (1) applying class-specific priors estimated from training data (dubbed “proj. w/ class prior”), and (2) training a model to predict the pose&shape for the 2D box of an object (dubbed “proj. w/ learning”). Table 1 provides their results, showing that both methods underperform our Pro3D. This demonstrates the significant role of Pro3D’s 3D detection backbone, which learns to better lift 2D detections and estimate objects’ pose and shape.
One may ask whether Pro3D’s performance gains stem from using a powerful 2D detector or its lifting method. First, using powerful 2D detectors does help. We can compare FCOS3D which adds a 3D head on top of a 2D detector for 3D detection. As the original FCOS3D exploits a Res101 backbone, we improve it by using the stronger detector DINO (with a Transformer backbone) and finetune it on the DAIR-V2X-I dataset. We name this improved version FCOS3D+. Table 1 shows that FCOS3D+ indeed improves over FCOS3D. Moreover, Fig. 7 justifies this further: using stronger 2D detectors improves 3D detection further. Second, the lifting method matters, as shown from the comparison between FCOS3D+ and our Pro3D, both of which use DINO as the 2D detector. Pro3D significantly outperforms FCOS3D+, owing to Pro3D’s 3D backbone dedicated to lifting 2D detections. Table 1 also shows that stronger lifting methods (BEVSpread>>BEVHeight>>BEVDepth) yield more gains.
@@ 4.3 Ablation Study and Analysis
Exploiting a scene prior improves 3D detection. We concatenate a testing image and the scene prior (a “blank” RGB background image as shonw in Fig. 4) as a 6-channel input to the 3D detector. Table 3 shows that adding the scene prior remarkably improves roadside 3D detection. To study how different scene prior affects the final performance, we generate scene priors over training frames captured at night. Table 4 shows that our method is quite robust to scene priors that have different illuminations.
Exploiting a scene prior particularly improves small object detection. Recall that the scene prior is a background reference. It helps precise 2D localization of small objects by grounding them at their contacts on the ground. This helps 3D detectors lift the precise 2D boxes for better 3D detections. Fig. 6 demonstrates this by comparing Pro3D’s predictions without versus with such a scene prior.
Figure 7: We train different 2D detectors (YOLOV7 [wang2022yolov7] and DINO [zhang2022dino]) on different amounts of 2D annotations (provided by nuImages [caesar2020nuscenes]) for Pro3D. We measure the 3D detection by averaging AP over the three classes vehicle, cyclist and pedestrian. Training on more 2D annotations leads to better 2D detectors and importantly, increases 3D detection performance of Pro3D.
Exploiting a stronger 2D detection yields better 3D detection performance. We train stronger 2D detectors (e.g., YOLOV7 [wang2022yolov7] and DINO [zhang2022dino]) on more 2D annotations from nuImages [caesar2020nuscenes]. Fig. 7 plots 3D detection performance by our Pro3D versus 2D detection performance (by training different 2D detectors). Results show that exploiting stronger 2D detectors greatly improves 3D detection performance with our Pro3D.
Table 5: Comparison of different prompting methods. As a baseline, we train BEVHeight with scene prior (Table 3). First, all prompting methods remarkably improves over BEVHeight. Second, prompting with predicted 2D box coordinates (“fuse 2D centers”) outperforms features. Third, additionally adding the predicted labels in prompts improves further.
@@
BEVHeight w/ scene prior 79.2 67.7 67.8 62.5 62.3 62.7 44.3 42.5 42.7
@@
What to use as prompts? We study what information from the 2D detector can be used as prompts to facilitate 3D detector training, including the intermediate feature map, the feature vectors located at a 2D detection, and even the simplistic 2D detection output (i.e., predicted box coordinates and class label). They require different fusion methods to accommodate their different dimensions. For example, when using feature maps, one can concatenate them with the feature map of 3D detector backbone, while using feature vectors or detections require fusing vectors with the feature maps. Hence, we use concatenation to fuse feature maps, and an attention module to fuse the feature vectors. Table 5 compares them. Surprisingly, using the simple detections (2D boxes and class labels) as prompts performs the best among the three methods: it achieves [68.5, 66.2, 47.2] AP for Hard examples of the three classes [vehicle, cyclist, pedestrian]. Furthermore, additionally incorporating the predicted class labels in prompts boosts the performance to [68.9, 68.2, 50.1] AP on Hard examples of the three classes! In hindsight, fusing box coordinates and labels provides both semantic and spatial information for the 3D detector to leverage, facilitating 3D detection. In contrast, fusing feature maps does not pinpoint the location of the objects and hence does not help 3D detector training.
Table 6: Stage-wise training (SW) vs. joint training (JT) for Pro3D. We report AP metrics at IoU=0.5, 0.25, 0.25 in BEV on the DAIR-V2X-I benchmark. As a baseline, we train Pro3D with BEVHeight and scene prior. Stage-wise training allows us to train separate 2D and 3D detector backbones, i.e., their backbones are not shared (NS). This allows us to exploit more 2D annotated data (MD) to train better 2D detectors, e.g., using nuImages [caesar2020nuscenes] to pretrain a 2D detector. Results convincingly show that SW outperforms JT for roadside monocular 3D detection. With NS and MD, Pro3D yields more notable improvements.
@@
Stage-wise training vs. joint training. We test different training paradigms in our Pro3D, e.g., (1) joint end-to-end training the 3D detector using both 2D and 3D losses, and (2) stage-wise training the 2D detector first and then the 3D detector. Note that stage-wise training allows us to train separate 2D and 3D detector backbones, i.e., their backbones are not shared (NS). This further makes it convenient to exploit more 2D annotated data (MD) to train better 2D detectors, e.g., using nuImages [caesar2020nuscenes] to pretrain a 2D detector. Results in Table 6 show that stage-wise training outperforms joint training for roadside monocular 3D detection.
How to train 3D detectors with 2D detection prompts? To train the 3D detector, we can either use predicted 2D boxes or the ground-truth as prompts. Table 7 compares their performance w.r.t AP in BEV on the DAIR-V2X-I benchmark, showing that the adpoting the former yields better performance.
@@
w/ grnd 82.7 70.1 70.3 70.5 70.4 70.6 52.8 51.7 51.3
w/ pred 82.9 70.7 70.8 71.3 71.2 71.4 53.9 52.9 52.6
Table 8: Comparison of inference time and parameters between Pro3D and BEVHeight. AP means the average AP in BEV over all the three classes in DAIR-V2X-I; “#param” counts the parameters (in million); time is the wall-clock in ms, FPS is frames per second. We use two different 2D detectors in our Pro3D, namely YOLOV7 and DINO. We list the number of parameters of 3D detector plus 2D detector. Our Pro3D “w/ YOLOV7” boosts 3D detection performance by 7.9 AP over the prior art BEVHeight, with faster inference speed (15.4 vs. 13.0 FPS) and fewer parameters (86.6M vs. 94.7M)!
@@
Latency, parameters, and training time. Recall that Pro3D achieves the best performance by using separate 2D and 3D detectors without sharing the visual encoder (Table 6). We study this model in Table 8 w.r.t model parameters, the wall-clock time, and FPS in processing a single RGB frame, as well as training time. For inference, we run the 2D detector (either YOLOV7 or DINO) and the 3D detector encoder in parallel. It is worth noting that Pro3D does not require an FPN for proposal detection as the 2D detector provides pinpointed detections already; it also replaces Res18 encoder (as adopted in prior methods [yang2023BEVHeight]) with a single convolution layer (Sec. 4.1). These changes make our Pro3D faster than the original BEVHeight in inference, regardless exploiting either YOLOV7 [wang2022yolov7] or DINO [yang2023BEVHeight] as the 2D detector. The changes also accelerate training. Table 9 compares the training times (in minutes) of BEVHeight and our Pro3D (built on BEVHeight). One might have thought that Pro3D requires more training time as it finetunes a 2D detector (YOLOv7 or DINO); in contrast, it needs less training time: a finetuned 2D detector provides a better base model for 3D training and thus reduces the total epochs needed. Further, owing to the modification on BEVHeight for Pro3D (removing FPN, replacing MLP head with a single convolution layer), Pro3D requires less training time in each epoch. In summary, Pro3D not only achieves better performance than BEVHeight, but also has faster inference and training speed.
Table 9: Comparison of training time between Pro3D and BEVHeight. Following Table 8, we build Pro3D with BEVHeight for fair comparison. One might think, as Pro3D finetunes 2D detectors (YOLOv7 or DINO), it should need more training time. Quite differently, a finetuned 2D detector provides a better base model for 3D training and thus reduces the total epochs needed. As Pro3D does not need an FPN and replaces the BEVHeight’s Res18-based BEV encoder with a single convolution layer, it needs less time in each epoch.
Method 2D det. ft. min/epoch #epochs total time in minutes
BEVHeight n/a 13 75 $13\times 75=975$
Pro3D + YOLOV7 35 11 40 $35+11\times 40=475$
Pro3D + DINO 60 11 40 $60+11\times 40=500$